# SN 2021foa: Deriving a continuity between SN IIn and SN Ibn

### v1

@@ Abstract
We present the long-term photometric and spectroscopic monitoring campaign of a transitioning SN IIn/Ibn from 10.8 d to 150.7 d post -band maximum. SN 2021foa shows prominent He i lines which are comparable in strength to the H line around peak luminosity, placing SN 2021foa between the SN IIn and SN Ibn populations. The spectral comparison with SNe IIn and SNe Ibn shows that it resembles the SN IIn population pre-maximum, becomes intermediate between SNe IIn/Ibn and at post-maximum matches with SN IIn 1996al. The photometric evolution shows a precursor at 50 d and a light curve shoulder around 17 d, which matches well with the light curve of the Type IIn SN 2016jbu and some other SNe with interactions. The peak luminosity and color evolution of SN 2021foa are consistent with those of most SNe IIn and SNe Ibn in our comparison sample. SN 2021foa shows the unique case of a SN IIn where the P-Cygni features in H appear at later stages, arising either due to complex geometry of the CSM or an interaction of the ejecta with a CSM shell/disk (similar to SNe 2009ip and 2015bh). Temporal evolution of the H profile favours a disk-like CSM geometry (CSM having both H and He) with a narrow (500 – 1200 km s-1) component, intermediate width (3000 – 8000 km s-1) and broad component in absorption early on. Hydrodynamical lightcurve modelling can be well-reproduced by a two-component CSM structure with different densities ( r-2 – r-5), mass-loss rates (10-3 – 10-1 M⊙ yr-1) assuming a wind velocity of 1000 km s-1 and having a CSM mass of 0.18 M⊙. The overall lightcurve and spectral evolution indicates that SN 2021foa most likely originated from a LBV star transitioning to a WR star with the mass-loss rate increasing in the period from 5 to 0.5 years before the explosion or it could be due to a binary interaction.
@@
spectroscopy, photometry, supernovae (SNe), supernova (SN), SN 2021foa1 Introduction
Massive stars that eventually undergo core collapse when surrounded by some dense circumstellar material (CSM) are known as Type IIn/Ibn supernovae (SNe) (Schlegel, 1990; Filippenko, 1997; Fraser, 2020). This is signified in spectra by a bright, blue continuum and narrow emission lines at early times. SNe of Type IIn display H-emission lines with multi-component profiles showing narrow-width lines (NW), Intermediate-width (IW) lines and Broad width lines (BW). Narrow ( 100 500 km s-1) components arise mostly in the photo-ionized, slow-moving CSM. Intermediate width emission lines ( 1000 km s-1) arise from either electron scattering of photons in narrower lines or emission from gas shocked by supernova (SN) ejecta. Some events also show very broad emission or absorption features ( 10,000 km s-1) arising from fast ejecta, typically associated with material ejected in the core-collapse explosion. An interesting extension of the Type IIn phenomenon was illustrated by SN 2006jc (Foley et al., 2007; Pastorello et al., 2008b), which exhibited narrow He i lines instead of NW/IW Hydrogen lines as spectral signatures of strong CSM interaction. Spectra of SN 2006jc had intermediate-width lines similar to those of SNe IIn (2000 – 3000 km s-1), but seen mainly in He i emission lines – there was only a trace amount of Hydrogen in the spectra. It is therefore referred to as a ‘SN Ibn’ event, instead of a SN IIn. It has important implications for understanding the broader class of SNe IIn and Ibn with CSM interaction because it is also one of the few Type Ibn SN that was observed to have a non-terminal LBV-like outburst just 2 years prior to explosion (Itagaki et al., 2006; Pastorello et al., 2007). Other than the traditional definition, the existence of transitional events that change type between Type IIn and Type Ibn over time (e.g Pastorello et al., 2015a; Reguitti et al., 2022) suggests a continuum in the CSM properties of these events and consequently in the mass-loss history of their progenitor stars. We now also have a couple of new candidates of an interesting class of events named SNe Icn, which show narrow narrow emission lines of C and O, but are devoid of H and He (Gal-Yam et al., 2022; Perley et al., 2022; Fraser et al., 2021; Pellegrino et al., 2022; Davis et al., 2023; Nagao et al., 2023). Pursiainen et al. (2023) shows that the appearance of prominent narrow He emission lines in the spectra of SN 2023emq (Icn), around maximum light are typical of a SN Ibn. The family of interacting SNe thus occupy a unique space, but probably linked by a continuum of outer envelopes.
@@
et al., 2020), an inflation of the progenitors radius that triggers violent binary interactions like collisions or mergers before the core-collapse event (Smith &
Arnett, 2014), or just the expansion of the envelope pre-collapse in massive stars (Mcley &
Soker, 2014).
@@
The traditional LBV stars as progenitors of Type IIn SNe are generally bright, blue and varying (Weis & Bomans, 2020). Smith et al. (2012b); Pastorello et al. (2015a, 2008b) introduced the first class of these events SNe 2005la and 2011hw, which as per the interpretation has progenitor star exploded as core-collapse while transitioning from the LBV to WR star phase. Theoretical models show that the event rate of appropriate binary mergers may match the rate of SNe with immediate LBV progenitors; and that the progenitor birthrate is 1 of the CCSNe rate (Justham et al., 2014). Observationally, the LBV/SN IIn connection inferred from properties of the CSM is reinforced by the detection of luminous LBV-like progenitors of three SNe IIn (Gal-Yam & Leonard, 2009; Kochanek et al., 2011; Smith et al., 2011). The standard evolution models instead suggest that massive stars are supposed to undergo only a very brief (104 – 105 yr) transitional LBV phase, and then spend 0.5-–1 Myr in the core-He burning Wolf-Rayet (WR) phase before exploding as a stripped envelope SN Ib/Ic (Heger et al., 2003). This discrepancy between observational and theoretical numbers most likely exists due to insufficient mass-loss rate estimates, not considering binary evolution scenario and also not taking into account the criticality of the LBV phase (Smith et al., 2012b). This accounts for the fact that stellar evolution models are missing essential aspects of the end stages of massive stars.
@@
While SNe IIn explosions make up 8–9 percent of all core-collapse SNe in the Lick Observatory Supernova Search sample (Li et al., 2011), the Type Ibn events like SN 2006jc represent a substantially smaller fraction. The fraction of SN 2006jc like event in this case constituted only 1 percent of the core-collapse sample which agrees with an independent estimate of the fraction of SN Ibn events by Pastorello et al. (2008a). Perley et al. (2020) updated this fraction for the Zwicky Transient Factory and found that SNe IIn consitute 14.2 percent of H-rich CCSNe while SNe Ibn constitute 9.2 percent of H-poor CCSNe. Given their rare occurrence, additional examples are valuable to demonstrate the diversity of the subclass. Among this whole sample, there are very few members of this peculiar SN class like SNe 2005la and 2011hw which had prominent narrow H and He lines (Pastorello et al., 2008b). This motivates us to study another rare case of transitioning Type IIn/Ibn SNe, which belongs to the same category as SNe 2005la and 2011hw.
@@
SN 2021foa were already investigated by Reguitti et al. 2022 (hereafter R22). Here we present further detailed photometric and spectroscopic observations of SN 2021foa, which exhibits both H and He emission features and shows similarities with both Type IIn SNe and Type Ibn SNe at distinct phases of its evolution. SN 2021foa like SNe show similarities in photometric evolution with both Type IIn and Type Ibn SNe (Hosseinzadeh et al., 2017; Ransome et al., 2021), but their diverse spectroscopic behaviour needs to be explored further to understand the division.
SN 2021foa (a.k.a. ASASSN-21dg, ATLAS21htp, PS21cae) was discovered by the All Sky Automated Survey for SuperNovae (ASAS-SN; Shappee et al., 2014) on 15 March 2021 (MJD=59288.45) at a Sloan apparent magnitude of 15.9, with the last non-detection 10 days earlier, at of 17.9 mag (Stanek & Kochanek, 2021). The SN was discovered at a RA and Dec of =13:17:12.29, =17:15:24.19 (J2000) in a barred spiral galaxy IC 863 at a redshift of z=0.008386 (Pisano et al., 2011). The observational and data reduction details of this SN are presented Section 2. The spectroscopic classification was done on 09-04-2021 using Nordic Optical Telescope (NOT) (Angus, 2021). The spectroscopic features and the details on the spectral luminosity estimates are provided in Section 3. We also show a comparison with other spectra of SNe IIn and SNe Ibn to highlight similarities and dissimilarities in Section 3.2. We discuss the photometric evolution, color, and absolute magnitude of SN 2021foa along with other members of the comparison sample in Section 4. The detailed lightcurve and spectral modelling is discussed in Section 5. Section 6 gives an estimate of mass-loss rates in these SNe and relates it with progenitor activity. Section 7 discusses the overall scenario of SN 2021foa and we summarise our results in Section 8.
@@
This work presents an extended analysis on SN 2021foa after R22. In their analysis, R22 showed that SN 2021foa belongs to sub-class of SN IIn which are labelled as SN IId (Benetti et al., 2016) and shows prominent narrow H early on with ejecta signatures later on. SN 2021foa, however, showed early prominent signatures of He i 5876 Å than other SNe IIn like 2009ip, 2016jbu. R22 quoted that SN 2021foa may be part of a bridge connecting H-rich SN 2009ip-like and Type Ibn SN events, indicating the possible existence of a continuum in properties, mass-loss history and progenitor types between these two types of peculiar transients. In our paper, we did a more robust modelling of the lightcurve, spectra and derived the physical parameters associated with the explosion. R22 indicated that SNe IId are probably connected to those objects by having similar a progenitor with LBVs transitioning to WR phase, but with a different mass-loss history or observed with a different orientation. We indeed, notice that our estimated mass-loss rates changes at different phases in the evolution of the SN, and spectral modelling shows an asymmetric CSM structure giving rise to H and He i at different strengths. Thus, our results are in concordance with what has been predicted by R22 and also an more elaborate description of it.
@@ 2 Observations and Data Reduction
2.1 Optical Observations
We observed SN 2021foa in UBgVriocRIJHK bands from day 34.3 to 150 d post V-band maximum (see section 2.3). The oc-band ATLAS data was reduced and calibrated using the techniques mentioned in (Tonry et al., 2018). The imaging observations were carried out using the 1.5m Kanata telescope (KT; Kawabata et al., 2008) of Hiroshima University; Japan, 3.8m Seimei Telescope (Kurita et al., 2020) of Kyoto University at Okayama observatory, Japan. Several bias, dark, and twilight flat frames were obtained during the observing runs along with science frames. For the initial pre-processing, several steps, such as bias-subtraction, flat-fielding correction, and cosmic ray removal, were applied to raw images of the SN. We used the standard tasks available in the data reduction software IRAF111IRAF stands for Image Reduction and Analysis Facility distributed by the National Optical Astronomy Observatory, operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. for carrying out the pre-processing. Multiple frames were taken on some nights and co-added in respective bands after the geometric alignment of the images to increase the signal-to-noise ratio.
@@ 2.1 Optical Observations
Low-resolution (R ) optical spectroscopic observations were carried out using the FLOYDS spectrographs mounted on the LCO 2m telescopes. The 1D wavelength and flux calibrated spectra were extracted using the floydsspec333https://github.com/LCOGT/floyds_pipeline pipeline (Valenti et al., 2014). Spectroscopic observations were also carried out using the KOOLS-IFU (Matsubayashi et al., 2019) on the Seimei Telescope. Our spectral coverage spans from 10.8 d to +69.5 d. The spectra with KOOLS-IFU were taken through optical fibers and the VPH-blue grism. The data reduction was performed using the Hydra package in IRAF (Barden, 1994) and a reduction software developed for KOOLS-IFU data444http://www.o.kwasan.kyoto-u.ac.jp/inst/p-kools. For each frame, we performed sky subtraction using a sky spectrum created by combining fibers to which the contributions from the object are negligible. Arc lamps of Hg, Ne, and Xe were used for wavelength calibration. Finally, the spectra were corrected for the heliocentric redshift of the host galaxy. The slit loss corrections were done by scaling the spectra with respect to the SN photometry. The log of spectroscopic observations is reported in Table 4.
Along with that, we also obtained high-resolution spectroscopic data with High Dispersion Spectrograph (HDS) mounted on Subaru Telescope on 2021 April 22 (UT). The echelle setup was chosen to cover the wavelength range of 5700–7100 Å with a spectral resolution of 50,000 in the Red Cross Disperser mode. We followed standard procedures to reduce the data. The wavelength calibration was performed using Th–Ar lamps. A heliocentric velocity correction was applied to each spectrum. The sky subtraction was performed using data at an off-target position in the target frames. The spectra were not flux calibrated by a standard star but were scaled to photometric fluxes at similar epochs to account for any flux losses.
@@ 2.2 Radio Observations
The observing campaign of SN 2021foa was carried out using the Giant Meterwave Radio Telescope (GMRT), Pune, India in Bands 4 and 5 (PI: Poonam Chandra). There was no detection of the source on observations dated 11 January 2022 and 31 March 2022. Our rms obtained in Band 5 (1.265 GHz) and Band 4 (0.745 GHz) are 33 and 215 Jy and the 3-sigma limits corresponding to the non-detections are 100 Jy and 645 Jy in the two bands. There was a nearby radio-bright galaxy with extended emission which was contaminating the SN location due to which our Band 4 rms are high. We also checked the Very-Large Array Telescope archive and did not find any detection of this source. The obtained radio luminosities at these phases for Band 5 and Band 4 are 1.455 1026 erg sec-1 and 9.385 1026 erg sec-1. Figure 18 of Gangopadhyay et al. (2023) shows the radio luminosity of a group of core-collapse SNe which includes SNe IIn. We see that the minimum radio luminosity driving SNe IIn are in between 1027 - 1029 erg sec-1. We do expect a radio emission at a phase of 584 d (1st observation) if it were a SNe IIn. However, given the fast decaying light curve of SN 2021foa, it’s expected that the radio power will decrease, and thus, we get only radio upper limits.

### v3

# SN 2021foa: the bridge between SN IIn and Ibn

@@ Abstract
We present the long-term photometric and spectroscopic analysis of a transitioning SN IIn/Ibn from 10.8 d to 150.7 d post -band maximum. SN 2021foa shows prominent He i lines comparable in strength to the H line around peak, placing SN 2021foa between the SN IIn and SN Ibn populations. The spectral comparison shows that it resembles the SN IIn population at pre-maximum, becomes intermediate between SNe IIn/Ibn and at post-maximum matches with SN IIn 1996al. The photometric evolution shows a precursor at 50 d and a light curve shoulder around 17d. The peak luminosity and color evolution of SN 2021foa are consistent with most SNe IIn and Ibn in our comparison sample. SN 2021foa shows the unique case of a SN IIn where the narrow P-Cygni in H becomes prominent at 7.2 days. The H profile consists of a narrow (500 – 1200 km s-1) component, intermediate width (3000 – 8000 km s-1) and broad component in absorption. Temporal evolution of the H profile favours a disk-like CSM geometry. Hydrodynamical modelling of the lightcurve well reproduces a two-component CSM structure with different densities ( r-2 – r-5), mass-loss rates (10-3 – 10-1 M⊙ yr-1) assuming a wind velocity of 1000 km s-1 and having a CSM mass of 0.18 M⊙. The overall evolution indicates that SN 2021foa most likely originated from a LBV star transitioning to a WR star with the mass-loss rate increasing in the period from 5 to 0.5 years before the explosion or it could be due to a binary interaction.
@@
spectroscopy, photometry, supernovae (SNe)1 Introduction
Massive stars that eventually undergo core collapse when surrounded by some dense circumstellar material (CSM) are known as Type IIn/Ibn supernovae (SNe) (Schlegel, 1990; Filippenko, 1997; Fraser, 2020). This is signified in spectra by a bright, blue continuum and narrow emission lines at early times. SNe of Type IIn display H-emission lines with multi-component profiles showing narrow-width lines (NW), Intermediate-width (IW) lines and Broad width lines (BW). Narrow ( 100 500 km s-1) components arise mostly in the photo-ionized, slow-moving CSM. Intermediate width emission lines ( 1000 km s-1) arise from either electron scattering of photons in narrower lines or emission from gas shocked by supernova (SN) ejecta. Some events also show very broad emission or absorption features ( 10,000 km s-1) arising from fast ejecta, typically associated with material ejected in the core-collapse explosion. An interesting extension of the Type IIn phenomenon was illustrated by SN 2006jc (Foley et al., 2007; Pastorello et al., 2008b), which exhibited narrow He i lines instead of NW/IW Hydrogen lines as spectral signatures of strong CSM interaction. Spectra of SN 2006jc had intermediate-width lines similar to those of SNe IIn (2000 – 3000 km s-1), but seen mainly in He i emission lines – there was only a trace amount of Hydrogen in the spectra. SN 2006jc is therefore referred to as a ‘SN Ibn’ event, instead of a SN IIn. It has important implications for understanding the broader class of SNe IIn and Ibn with CSM interaction because it is also one of the few Type Ibn SN that was observed to have a non-terminal LBV-like outburst just 2 years prior to explosion (Itagaki et al., 2006; Pastorello et al., 2007). Other than the traditional definition, the existence of transitional events that change type between Type IIn and Type Ibn over time (e.g Pastorello et al., 2015a; Reguitti et al., 2022) suggests a continuum in the CSM properties of these events and consequently in the mass-loss history of their progenitor stars. We now also have a couple of new candidates of an interesting class of events named SNe Icn, which show narrow narrow emission lines of C and O, but are devoid of H and He (Gal-Yam et al., 2022; Perley et al., 2022; Fraser et al., 2021; Pellegrino et al., 2022; Davis et al., 2023; Nagao et al., 2023). Pursiainen et al. (2023) shows that the appearance of prominent narrow He emission lines in the spectra of SN 2023emq (Icn), around maximum light are typical of a SN Ibn. The family of interacting SNe thus occupy a unique space, but probably linked by a continuum of outer envelopes.
@@
et al., 2020), an inflation of the progenitors radius that triggers violent binary interactions like collisions or mergers before the core-collapse event (Soker, 2013; Smith &
Arnett, 2014; Mcley &
Soker, 2014a), or just the expansion of the envelope pre-collapse in massive stars (Mcley &
Soker, 2014b).
@@
The traditional LBV stars as progenitors of Type IIn SNe are generally bright, blue and varying (Weis & Bomans, 2020). Smith et al. (2012b); Pastorello et al. (2015a, 2008b) introduced the first class of these events SNe 2005la and 2011hw, which as per the interpretation has progenitor star exploded as core-collapse while transitioning from the LBV to WR star phase. Theoretical models show that the event rate of appropriate binary mergers may match the rate of SNe with immediate LBV progenitors; and that the progenitor birthrate is 1 of the CCSNe rate (Justham et al., 2014). Observationally, the LBV/SN IIn connection inferred from properties of the CSM is reinforced by the detection of luminous LBV-like progenitors of three SNe IIn (Gal-Yam & Leonard, 2009; Kochanek et al., 2011; Smith et al., 2011). Standard evolution models instead suggest that massive stars are supposed to undergo only a very brief (104 – 105 yr) transitional LBV phase and then spend 0.5-–1 Myr in the core-He burning Wolf-Rayet (WR) phase before exploding as a stripped envelope SN Ib/Ic (Heger et al., 2003). This discrepancy between observational and theoretical numbers most likely exists due to insufficient mass-loss rate estimates, not considering binary evolution scenario and also not taking into account the criticality of the LBV phase (Smith et al., 2012b). This accounts for the fact that stellar evolution models are missing essential aspects of the end stages of massive stars.
@@
While SNe IIn explosions make up 8–9 percent of all core-collapse SNe in the Lick Observatory Supernova Search sample (Li et al., 2011; Perley et al., 2020), the Type Ibn events like SN 2006jc represent a substantially smaller fraction. The fraction of SN 2006jc like event in this case constituted only 1 percent of the core-collapse sample which agrees with an independent estimate of the fraction of SN Ibn events by Pastorello et al. (2008a). Perley et al. (2020) updated this fraction for the Zwicky Transient Factory and found that SNe IIn consitute 14.2 percent of H-rich CCSNe while SNe Ibn constitute 9.2 percent of H-poor CCSNe. Given their rare occurrence, additional examples are valuable to demonstrate the diversity of the subclass. Among this whole sample, there are very few members of this peculiar SN class like SNe 2005la and 2011hw which had prominent narrow H and He lines (Pastorello et al., 2008b; Kool et al., 2021; Farias, 2024). This motivates us to study another rare case of transitioning Type IIn/Ibn SNe, which belongs to the same category as SNe 2005la and 2011hw.
@@
SN 2021foa were already investigated by Reguitti et al. 2022 (hereafter R22). Here we present further detailed photometric and spectroscopic observations of SN 2021foa, which exhibits both H and He emission features and shows similarities with both Type IIn and Type Ibn SNe at distinct phases of its evolution. SN 2021foa show similarities in photometric evolution with both Type IIn and Type Ibn SNe (Hosseinzadeh et al., 2017; Ransome et al., 2021), but their diverse spectroscopic behaviour needs to be explored further to understand the division.
@@
This work presents an extended analysis on SN 2021foa after R22. In their analysis, R22 showed that SN 2021foa belongs to sub-class of SN IIn which are labelled as SN IId (Benetti et al., 2016) and shows prominent narrow H early on with ejecta signatures later on. SN 2021foa, however, showed early prominent signatures of He i 5876 Å than other SNe IIn like 2009ip, 2016jbu. R22 quoted that SN 2021foa may be part of a bridge connecting H-rich SN 2009ip-like and Type Ibn SNe, indicating the possible existence of a continuum in properties, mass-loss history and progenitor types between these two types of peculiar transients. In our paper, we did a more robust modelling of the lightcurve, spectra and derived the physical parameters associated with the explosion. R22 indicated that SNe IId are probably connected to those objects by having similar a progenitor with LBVs transitioning to WR phase, but with a different mass-loss history or observed with a different orientation. We indeed, notice that our estimated mass-loss rates changes at different phases in the evolution of the SN, and spectral modelling shows an asymmetric CSM structure giving rise to H and He i at different strengths. Thus, our results are in concordance with what has been predicted by R22 and also an more elaborate description of it.
@@ 2 Observations and Data Reduction
2.1 Optical Near-Infrared Observations
We observed SN 2021foa in UBgVriocRIJHK bands from day 34.3 to 150 d post V-band maximum (see section 2.3). The oc-band ATLAS data was reduced and calibrated using the techniques mentioned in (Tonry et al., 2018). Imaging observations in BVRIJHK were carried out using the 1.5m Kanata telescope (KT; Kawabata et al., 2008) of Hiroshima University; Japan. Several bias, dark, and twilight flat frames were obtained during the observing runs along with science frames. For the initial pre-processing, several steps, such as bias-subtraction, flat-fielding correction, and cosmic ray removal, were applied to raw images of the SN. We used the standard tasks available in the data reduction software IRAF111IRAF stands for Image Reduction and Analysis Facility distributed by the National Optical Astronomy Observatory, operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. for carrying out the pre-processing. Multiple frames were taken on some nights and co-added in respective bands after the geometric alignment of the images to increase the signal-to-noise ratio.
@@ 2.1 Optical Near-Infrared Observations
Low-resolution (R ) optical spectroscopic observations were carried out using the FLOYDS spectrographs mounted on the LCO 2m telescopes. The 1D wavelength and flux calibrated spectra were extracted using the floydsspec333https://github.com/LCOGT/floyds_pipeline pipeline (Valenti et al., 2014). Spectroscopic observations were also carried out using the KOOLS-IFU (Matsubayashi et al., 2019) on the Seimei Telescope. Our spectral coverage spans from 10.8 d to +69.5 d. The spectra with KOOLS-IFU were taken through optical fibers and the VPH-blue grism. The data reduction was performed using the Hydra package in IRAF (Barden, 1994) and a reduction software developed for KOOLS-IFU data444http://www.o.kwasan.kyoto-u.ac.jp/inst/p-kools (Proposal numbers: 21A-N-CT02, 21A-O-0004, 2021-04-05 21A-K-CT02, 2021-04-14 21A-N-CT02, 2021-05-03 21A-O-0008). For each frame, we performed sky subtraction using a sky spectrum created by combining fibers to which the contributions from the object are negligible. Arc lamps of Hg, Ne, and Xe were used for wavelength calibration. Finally, the spectra were corrected for the heliocentric redshift of the host galaxy (see for reference Section 3). The slit loss corrections were done by scaling the spectra with respect to the SN photometry. The log of spectroscopic observations is reported in Table 4.
Along with that, we also obtained high-resolution spectroscopic data with High Dispersion Spectrograph (HDS) mounted on Subaru Telescope on 2021 April 22 (UT). This observation was done as a part of the Subaru Proposal No S21A-014 (PI: Keiichi Maeda). The Echelle setup was chosen to cover the wavelength range of 5700–7100 Å with a spectral resolution of 50,000 in the Red Cross Disperser mode. We followed standard procedures to reduce the data. The wavelength calibration was performed using Th–Ar lamps. A heliocentric velocity correction was applied to each spectrum. The sky subtraction was performed using data at an off-target position in the target frames. The spectra were not flux calibrated by a standard star but were scaled to photometric fluxes at similar epochs to account for any flux losses.
@@ 2.2 Radio Observations
The observing campaign of SN 2021foa was carried out using the Giant Meterwave Radio Telescope (GMRT), Pune, India in Bands 4 and 5 (PI: Poonam Chandra). There was no detection of the source on observations dated 11 January 2022 and 31 March 2022. Our rms obtained in Band 5 (1.265 GHz) and Band 4 (0.745 GHz) are 33 and 215 Jy and the 3-sigma limits corresponding to the non-detections are 100 Jy and 645 Jy in the two bands. There was a nearby radio-bright galaxy with extended emission which was contaminating the SN location due to which our Band 4 rms are high. We also checked the Very-Large Array Telescope archive and did not find any detection of this source. The obtained radio luminosities at these phases for Band 5 and Band 4 are 1.455 1026 erg sec-1 and 9.385 1026 erg sec-1. Figure 18 of Gangopadhyay et al. (2023) shows the radio luminosity of a group of core-collapse SNe which includes SNe IIn. We see that the minimum radio luminosity driving SNe IIn are in between 1027 - 1029 erg sec-1. We do expect a radio emission at a phase of 584 d (1st observation) if it were a SNe IIn. Previous cases of SNe 2006jd, 2010jl (Chandra et al., 2012; Chandra et al., 2015) have shown radio lightcurves peaking between 500 d - 700 d post explosion. Radio emission in SNe IIn is expected to be synchrotron emission, initially absorbed mainly by free–free absorption, while X-ray, optical emission is likely to have a thermal, radioactive origin. So, even if we expect the radio lightcurves to peak for SN 2021foa, given the fast decaying light curve of SN 2021foa, it’s expected that the radio power will decrease, and thus, we get only radio upper limits.